State the inverse for the function:
g(x) = (7/3)x + 2
I've never done a problem like this before and don't know where to start, I'd like to see the steps and the answer so I can understand what to do. Thank you!

Found 2 solutions by stanbon, funmath:
Answer by stanbon(75887)


g(x) = (7/3)x + 2
---------------
1st: Interchange x and y to get:
x = (7/3)y + 2
---------
2nd: Solve for y:
x-2 = (7/3)y
y = (3/7)(x-2)
---------
That is the inverse.
Cheers,
Stan H.

Answer by funmath(2933)

State the inverse for the function:
g(x) = (7/3)x + 2
:
1. Let g(x)=y
y=(7/3)x+2
:
2. Let x be y and y be x
x=(7/3)y+2
:
3. Solve for y
x-2=(7/3)x+2-2
x-2=(7/3)y
(3/7)(x-2)=(3/7)(7/3)y
(3/7)x-(6/7)=y
y=(3/7)x-(6/7)
:
4. Let y=g^(-1)(x)
g^(-1)(x)=(3/7)x-(6/7)
:
Happy Calculating!!!